OpenStudy (anonymous):

please help (INTEGRATION BY PARTS) ∫x^-2 ln(x) dv please show what u,du,dv,and v are equal to

zepdrix (zepdrix):

\[\Large\rm \int\limits x^{-2}\ln x~dx\]So when you see a natural log, one thing to think about is "Oh yah, the derivative of ln is a power of x... so maybe I can let that be my u since it will combine nicely with the other term after I differentiate it." I think that's what we'll want in this case.

zepdrix (zepdrix):

\[\Large\rm u=\ln x,\qquad\qquad dv=x^{-2}dx\]

zepdrix (zepdrix):

\[\Large\rm du=\frac{1}{x}dx,\qquad\quad v=-x^{-1}\]So we get something like that for our parts, yes? Make sense so far?

OpenStudy (anonymous):

yes cause we take the derivative of u and the antiderivative of dv right

zepdrix (zepdrix):

yes.

zepdrix (zepdrix):

I dunno why I did the different shapes underlines lol. But whatever, I did.